OpenStudy (anonymous):

Eliza and Jamie are both making cupcakes for a bake sale. Eliza needs 2 1/3cups of flour and Jamie needs 1 3/4 cups of flour. They have 4 cups of flour. Do they have enough for both recipes? I'll medal and fan @KamiBug @SuperflyTyGuy @bohotness

OpenStudy (superflytyguy):

Do you know how to convert them into improper fractions?

OpenStudy (anonymous):

no not really im not good with fractions

OpenStudy (superflytyguy):

okay. Here how we do it. 2 1/3 needs to be an improper fraction Multiply the denominator and add the numerator. 2 1/3 = 2 x 3 = 6 6 + 1 = 7 2 1/3 = 7/3

OpenStudy (bohotness):

oaky i got you love

OpenStudy (superflytyguy):

Same thing with 1 3/4 1 x 4 = 4 4 + 3 = 7 So 7/4

OpenStudy (bohotness):

:Dhope that helps

OpenStudy (superflytyguy):

Now we gotta find the LCD The denominators 3 and 4 have 12 in common

OpenStudy (superflytyguy):

7/3 = 28/12 7/4 = 21/12 Add both of those up 28/12 + 21/12 = 49/12 49/12 = 4 1/12 No, they DO NOT have enough flour